If you model the tube as an open-closed system, wh
Answer #1

for open-closed pipe:
fn = (2n+1)*v/4L

f1 = v/4*L
= 350/(4*1.3)
= 67 Hz

f2 = 3v/4*L
= 3*350/(4*1.3)
= 200 Hz

f3 = 5v/4*L
= 5*350/(4*1.3)
= 340 Hz